28.7 The magnetic field of 60 cm long solenoid increases by 0.075 T when the current increases by 10.7 A. What is the number of turns in the solenoid?
Answer #1

here,

the magnetic feild increase , dB = 0.075 T

length , l = 60 cm = 0.6 m

let the number of turns in the solenoid be N

curret , dI = 10.7 A

dB = u0 * N * dI /l

0.075 = 1.26 * 10^-6 * N * 10.7 /(0.6)

solving for N

N = 3.34 * 10^3 turns

the number of turns in the solenoid is 3.34 * 10^3 turns